y=-1/3x+2/3
y=7/3x+2/3

Answer by Boreal(15235)

I'm assuming this is to solve for x and y
y=-1/3x+2/3
y=7/3x+2/3
subtract the second from the first
0=-(8/3)x
x=0
y=2/3